OpenStudy (anonymous):

Why is it important to know how to convert from one form into another form? Help!

OpenStudy (anonymous):

@superdavesuper

OpenStudy (anonymous):

In mathematical forms please :)

OpenStudy (linn99123):

OHHHH

OpenStudy (linn99123):

*Hides face* idk D:

OpenStudy (anonymous):

Thats ok @linn99123 :)

OpenStudy (superdavesuper):

transform different math forms? like from y=mx+c to Ax+by+C=0? need more context on the Q itself.

OpenStudy (anonymous):

Well like in what situation and why would you need to convert a fraction to a decimal or vise versa.

OpenStudy (superdavesuper):

There are many different reasons and they are specific to the questions and forms. What is ur teach expecting u to write on this Q?

OpenStudy (anonymous):

Some thing about fractions and decimals, im sorry there is not much detail, but can you give one example?

OpenStudy (superdavesuper):

well like this: \[0.5 + \frac{ 1 }{ 2 }\] will be difficult to add. so u convert it into 0.5 + 0.5 = 1

OpenStudy (anonymous):

Thank you!! this is exactly what i needed! thank you so much you're the best!

OpenStudy (superdavesuper):

welcome...although im not sure thats what ur teach is expecting! good luck :)